# Localized Gaussians as Self-Attention Weights for Point Clouds Correspondence

### v1

@@ 2.1 Non-rigid shape matching
It is fundamental for many applications to map each point of a surface to one point of a second surface that has undergone a non-rigid deformation. For this problem, known as non-rigid shape matching, many contributions have been raised with a number of approaches. Descriptor-based methods long offered a solution by computing a vector of features, invariant to a number of transformations, for each point. Then, the matches are assigned according to a similarity score between points in the feature space [STDS14], [SY11], [TBW11]. A more comprehensive review of the matching methodologies can be found in [DYDZ22] and, in particular, for the methods that extrinsically solve the correspondence problem in [Sah20]. Thanks to the rise of machine learning, more data-driven methods have been proposed, and the transformer architecture has been proven to be a good solution for this problem. Some methods involving an encoder-decoder architecture, naturally designed to allow interactions between more inputs, have emerged; for example, in [TCM21], the Perceiver architecture [JGB21], originally proposed to target classification tasks across various modalities, is adapted to the shape matching task. Oppositely, in [RPM23], the authors proposed a shape matching solution with state-of-the-art competitive performances that exploits an encoder-only transformer architecture. This work pointed out the importance of introducing positional information about the points and conducted an extensive ablation study to identify the critical points of the architecture. In [RPM23], it has also been revealed that the attention patterns produced by the model are close to Gaussians centered around the points. Furthermore, it has been supposed that directly providing these patterns to the network could significantly speed up the learning procedure; in this work, we aim to explore this possibility, the advantages it brings and its drawbacks.
@@ 4 Methodology
To align with the dimensionality requirements of the Transformer encoder, the two-point clouds are projected into a $d$-dimensional space using a point-wise fully connected neural network. This transformation operates independently for each point, resulting in a higher-dimensional encoding of the original 3D coordinates. Instead, the output of the Transformer network is similarly projected back into a 3-dimensional space using a fully connected layer with learned weights.
The Transformer encoder architecture comprises multiple multi-head attention and feed-forward layers. Each multi-head attention layer consists of $h$ attention functions named heads. In our proposed architecture, some heads employ the conventional dot-product attention mechanism, while others utilize the Euclidean Gaussian function for attention weights. In both cases, residual attention is incorporated by adding the attention energy from the previous layer to the current layer.
@@
$\xi_{j,i}=softmax\left(\exp\left(\frac{\mathbf{E}^{2}}{2\sigma_{i}^{2}}\right)+\xi_{j-1,i}\right)$ (2)
where $\sigma_{i}$ is a fixed or learnable parameter of the network, and $\mathbf{E}$ is a $(2n+1)\times(2n+1)$ matrix and $\mathbf{E}_{p,q}$ is the Euclidean distance between the points $p$ and $q$ if they belong to the same shape, or 00 if the two points belong to different shapes.
@@
Matching and loss computation.
The output of the network is a matrix of shape $(2n+1)\times 3$, where each row corresponds to a point of the input mapped on the other shape. This matrix can be split into two point clouds, $\hat{X}$ and $\hat{Y}$, once removed the SEP element. $\hat{X}$ contains the points of $X$ remapped to fit the shape of $Y$, and conversely, $\hat{Y}$ contains the points of $Y$ remapped over $X$. Using the same setup described in [TCM21], the point-to-point correspondence is computed by solving the nearest neighbor assignment for each point $x\in X$ among the points of $\hat{Y}$, and likewise for each point $y\in Y$ over the points of $\hat{X}$. This correspondence allows to compute a simple loss as the sum of two losses $l_{X,Y}$ and $l_{Y,X}$ defined as:
$l=l_{X,Y}+l_{Y,X}=||\hat{Y}-X||_{2}^{2}+||\hat{X}-Y||_{2}^{2}$ (4)
@@ 5.1 Datasets and settings
To enhance the model’s robustness to rotations, for each input shape, we randomly apply one selected rotation from the following five possibilities: i) random rotation along all three axes within the interval $[0,2\pi]$; ii) random rotation along the x-axis within the interval $[0,2\pi]$; iii) random rotation along the y-axis within the interval $[0,2\pi]$; iv) random rotation along the z-axis within the interval $[0,2\pi]$; v) no rotation. Furthermore, to ensure the network is independent of the order of points in the point cloud and to prevent it from learning the order, a random permutation is applied to the input points in each shape before concatenating the two shapes. This permutation is then applied to the points before computing the losses to ensure the correct use of the ground truth correspondence.
@@
A configuration with no Gaussian attention heads, identified by the code 0gh600, is trained for 600 epochs with a batch size of 24. This configuration reproduces the APEAYN model trained with the same number of epochs, serving as a baseline for comparison.
@@
The models are tested on the shape matching task using the Faust1k dataset [BRLB14]. This dataset comprises 10 subjects in 10 different poses, represented by the same mesh, totaling 100 shapes with 1,000 points each. Additionally the models are tested on a noisy version of the same dataset to assess the robustness to noise of the methods.
@@
From the Faust1k dataset, 100 random pairings are selected as a test set. Each inference on a pair produces two matchings, $\pi_{X,Y}$ and $\pi_{Y,X}$, which map the points of $X$ to $Y$ and vice versa, respectively. Each matching computes the Chamfer loss against the ground truth, following the setup in [TCM21]. The matching that produces the lower Chamfer error is selected and used to compute the error for each point as the geodesic distance between the ground truth matching point and the closest point on the target shape to the inferred point. It is important to note that the network processes the mesh as a point cloud, retaining no information on connectivity. The geodesic distances are used solely as an evaluation metric.
The implementation of the training process, the evaluation and the models used for the experiments can be found at this repository: [The code will be released upon acceptance].
@@ 5.2 Results
Table 1: Comparison of the trained models results. For each model, we report the average geodesic error on the clean dataset and on the noisy dataset, the number of epochs of its learning, and the number of parameters. For the models we propose and for the baseline, we consider both the version obtained after 600 epochs and the one that reaches the best minimization of the loss (denoted as best.).
@@
It is possible to notice how the performances of the models with gaussian heads drastically decrease when faced with noisy data. The error on the noisy version of the Faust1k dataset is more than five times the error on the clean version for the 4gh600 and 4gh600.lis models, while it is only 65% more in the case of the 0gh600 model. This is due to the nature of the selected neighborhood, which is based solely on euclidean distances and is inherently very susceptible to noise. To mitigate this problem the 4gh600.lis.noise model has been introduced, it is trained identically to the 4gh600.lis model, with the addition of noise in the training set. The noise is injected on 50% of the points and extracted from a normal distribution $N(0,0.02)$, this additional data augmentation should also help in avoiding overfitting issues. The results, reported in Table 2 shows that the model is indeed more robust to noise with respect to the 4gh600.lis model but the performances on the clean dataset also decrease to the point it is slightly worse than the baseline, revealing that this is not the best solution to the problem and that further studies need to be conducted.
@@ 6 Analysis and Ablation
The model trained with the noisy dataset results in loss values higher than all the other models along all the epochs but the fist ones, where it registers loss values lower than the baseline. This is not unexpected as the loss values are computed on noisy data that is intrinsically more complex and challenging. It is however interesting to notice how the minimum loss epoch occurs earlier with respect to the other models and the plateau seems to be more pronunced. This suggests that the addition of noise in training helps the method adapt to noisy situations but also slows down the loss decrease, at least with the selected noise distribution.
@@ 6.2 Limitations
In this work, we focused on the shape matching task between human shapes only. Moreover, the tests were performed on a dataset that exclusively comprises point clouds of the same cardinality. Extensive testing on diverse and more challenging datasets, including both human and non-human shapes, is required to fully assess the applicability, generalizability, and potential drawbacks of the proposed methodology. The performaces on noisy data of the 4gh600.lis.noise show how the methodology could work but does require further studies to limit the matching deterioration on the clean dataset.

### v2

@@ 2.1 Non-rigid shape matching
It is fundamental for many applications to map each point of a surface to one point of a second surface that has undergone a non-rigid deformation. For this problem, known as non-rigid shape matching, many contributions have been raised with a number of approaches. Descriptor-based methods long offered a solution by computing a vector of features, invariant to a number of transformations, for each point. Then, the matches are assigned according to a similarity score between points in the feature space [STDS14], [SY11], [TBW11]. A more comprehensive review of the matching methodologies can be found in [DYDZ22] and, in particular, for the methods that extrinsically solve the correspondence problem in [Sah20]. While hand-crafted methods for shape correspondence produce good results, they present high developmental costs and are often domain-specific, making it difficult to apply them in other fields. Thanks to the rise of machine learning, more data-driven methods have been proposed, and the transformer architecture has been proven to be a good solution for this problem and to generalize well to a number of different contexts. Some methods involving an encoder-decoder architecture, naturally designed to allow interactions between more inputs, have emerged; for example, in [TCM21], the Perceiver architecture [JGB21], originally proposed to target classification tasks across various modalities, is adapted to the shape matching task. Oppositely, in [RPM23], the authors proposed a shape matching solution with state-of-the-art competitive performances that exploits an encoder-only transformer architecture. This work pointed out the importance of introducing positional information about the points and conducted an extensive ablation study to identify the critical points of the architecture. In [RPM23], it has also been revealed that the attention patterns produced by the model are close to Gaussians centered around the points. Furthermore, it has been supposed that directly providing these patterns to the network could significantly speed up the learning procedure; in this work, we aim to explore this possibility, the advantages it brings and its drawbacks.
@@ 4 Methodology
To align with the dimensionality requirements of the Transformer encoder, the two-point clouds are projected into a $d$-dimensional space using a point-wise fully connected neural network, with $d=512$ as set in the the APEAYN model. This transformation operates independently for each point, resulting in a higher-dimensional encoding of the original 3D coordinates. Instead, the output of the Transformer network is similarly projected back into a 3-dimensional space using a fully connected layer with learned weights.
The Transformer encoder architecture comprises multiple multi-head attention and feed-forward layers. Each multi-head attention layer consists of $h=8$ attention functions named heads. In our proposed architecture, some heads employ the conventional dot-product attention mechanism, while others utilize the Euclidean Gaussian function for attention weights. In both cases, residual attention is incorporated by adding the attention energy from the previous layer to the current layer.
The output of the network is a matrix of shape $(2n+1)\times 3$, where each row corresponds to a point of the input mapped on the other shape. This matrix can be split into two point clouds, $\hat{X}$ and $\hat{Y}$, once removed the SEP element. $\hat{X}$ contains the points of $X$ remapped to fit the shape of $Y$, and conversely, $\hat{Y}$ contains the points of $Y$ remapped over $X$.
@@
$\xi_{j,i}=softmax\left(\exp\left(-\frac{\mathbf{E}^{2}}{2\sigma_{i}^{2}}\right)+\xi_{j-1,i}\right)$ (2)
where $\sigma_{i}$ is a fixed or learnable parameter of the network, and $\mathbf{E}$ is a $(2n+1)\times(2n+1)$ matrix and $\mathbf{E}_{p,q}$ is the Euclidean distance between the points $p$ and $q$ if they belong to the same shape, or 0 if the two points belong to different shapes.
@@
Loss computation.
Given the two output point clouds $\hat{X}$ and $\hat{Y}$, that map the points of $X$ over the shape of $Y$ and vice versa, using the ground truth correspondence maps $\Pi_{X,Y}$ and $\Pi_{Y,X}$ we order the point clouds to respect the matching, meaning that $(x_{i},y_{i})\in\Pi_{X,Y}$ and $(y_{i},x_{i})\in\Pi_{Y,X}$ for any index $i$. This correspondence allows to compute a simple loss as the sum of two losses $l_{X,Y}$ and $l_{Y,X}$ defined as:
[…]
@@ 5.1 Datasets and settings
To enhance the model’s robustness to rotations, for each input shape, we randomly apply one selected rotation from the following five possibilities: i) random rotation along all three axes within the interval $0,2\pi]$; ii) random rotation along the x-axis within the interval $0,2\pi]$; iii) random rotation along the y-axis within the interval $0,2\pi]$; iv) random rotation along the z-axis within the interval $0,2\pi]$; v) no rotation. Furthermore, to ensure the network is independent of the order of points in the point cloud and to prevent it from learning the order, a random permutation is applied to the input points in each shape before concatenating the two shapes. This permutation is then applied to the points before computing the losses to ensure the correct use of the ground truth correspondence.
@@
A configuration with no Gaussian attention heads, identified by the code 0gh600, is trained for 600 epochs with a batch size of 24. This configuration reproduces the APEAYN model with a reduced number of epochs, from 5000 to 600, to have a fairer comparison to the other trained models.
@@
The models are tested on the shape matching task using the Faust1k dataset [BRLB14]. This dataset comprises 10 subjects in 10 different poses, represented by the same mesh, totaling 100 shapes with 1,000 points each. Additionally the models are tested on a noisy version of the same dataset, as done in [RPM23], to assess the robustness to noise of the methods. The noise on the test set is extracted from a normal distribution $N(0,0.01)$ and added on every point in the dataset.
@@
The evaluation of the results follows the setup in [TCM21, RPM23]. From the Faust1k dataset, 100 random pairings are selected as a test set. Each inference on a pair produces two output point clouds, $\hat{X}$ containing the points of $X$ remapped over the shape of $Y$, and $\hat{Y}$ remapping $Y$ over $X$. These constitute the data to compute two matchings, $\pi_{X,Y}$ and $\pi_{Y,X}$, which map the points of $X$ to $Y$ and vice versa, respectively. To select which matching is to be computed, for each inferred point cloud we compute the chamfer loss against the ground truth. The couple that produces the lower chamfer error is used to compute the matching. In particular, in the case $\hat{X}$ and $Y$ produce the lower loss value, for each point $x$ in the source point cloud $X$ we select the matching point $y_{x}$ in $Y$ as the point in $Y$ closest to $\hat{x}$, mapping of $x$ over the target shape $Y$. The error value is then defined as the geodesic distance between $y_{x}$ and $y$, point in $Y$ such that $(x,y)\in\Pi_{X,Y}$, ground truth matching. It is important to note that the network processes the mesh as a point cloud, retaining no information on connectivity. The geodesic distances are used solely as an evaluation metric.
The implementation of the training process, the evaluation and the models used for the experiments can be found at this repository: github.com/ariva00/GaussianAttention4Matching.
@@ 5.2 Results
Table 1: Comparison of the trained models results. For each model, we report the average geodesic error on the clean dataset (F1k error) and on the noisy dataset (F1knoise error), the number of epochs of its learning, and the number of parameters. For the models we propose and for the baseline, we consider both the version obtained after 600 epochs and the one that reaches the best minimization of the loss (denoted as best.).
@@
It is possible to notice how the performances of the models with gaussian heads drastically decrease when faced with noisy data. The error on the noisy version of the Faust1k dataset is more than five times the error on the clean version for the 4gh600 and 4gh600.lis models, while it is only 65% more in the case of the 0gh600 model. This is due to the nature of the selected neighborhood in the gaussian heads, which is based solely on euclidean distances and is inherently very susceptible to noise. To mitigate this problem the 4gh600.lis.noise model has been introduced, it is trained identically to the 4gh600.lis model, with the addition of noise in the training set. The noise is injected on 50% of the points and extracted from a normal distribution $N(0,0.02)$, this additional data augmentation should also help in avoiding overfitting issues. The results, reported in Table 2 shows that the model is indeed more robust to noise with respect to the 4gh600.lis model but the performances on the clean dataset also decrease to the point it is slightly worse than the baseline, revealing that this is not the best solution to the problem and that further studies need to be conducted.
@@ 6 Analysis and Ablation
The model trained with the noisy dataset results in loss values higher than all the other models along all the epochs but the first ones, where it registers loss values lower than the baseline. This is not unexpected as the loss values are computed on noisy data that is intrinsically more complex and challenging. It is however interesting to notice how the minimum loss epoch occurs earlier with respect to the other models and the plateau seems to be more pronunced. This suggests that the addition of noise in training helps the method adapt to noisy situations but also slows down the loss decrease, at least with the selected noise distribution.
@@ 6.2 Limitations
In this work, we focused on the shape matching task between human shapes only. Moreover, the tests were performed on a dataset that exclusively comprises point clouds of the same cardinality. With our work we aim to analyze in depth which type of geometric information enables a transformer architecture to achieve state-of-the-art results. For this reason, in our experiments, we limit our comparison to the most recent data-driven solution for shape matching based on transformers. Extensive testing on diverse and more challenging datasets, including both human and non-human shapes, is required to fully assess the applicability, generalizability, and potential drawbacks of the proposed methodology, as well as a broader comparison to state-of-the-art methods. The performaces on noisy data of the 4gh600.lis.noise show how the methodology could work but does require further studies to limit the matching deterioration on the clean dataset.
@@ 7 Conclusions
Acknowledgments
This work was partially supported by the MUR for REGAINS, the Department of Excellence DISCo at the University of Milano-Bicocca, the PRIN project GEOPRIDE Prot. 2022-NAZ-0115, CUP H53D23003400001, and by the NVIDIA Corporation with the RTX A5000 GPUs granted through the Academic Hardware Grant Program to the University of Milano-Bicocca for the project “Learned representations for implicit binary operations on real-world 2D-3D data”.